how can you simplify 2tanx + 2cotx anymore??
Answer by amalm06(224)

tan x=sin x/cos x
cot x=cos x/sin x
2(tan x)+2(cot x)=2(sin x/cos x)+2(cos x/sin x)
=2(sin x sin x+ cos x cos x)/(cos x sin x)

=2[sin^2x+cos^2x]/cos x sin x
=(2)(1)/cos x sin x
=2 sec x csc x (Answer)